OpenStudy (18jonea):

What is the mass of sulfur in 250 g of sulfuric acid, H2SO4? 16 g 0.33g 160 g 82 g @sweetburger

OpenStudy (18jonea):

@aaronq

OpenStudy (18jonea):

@Zale101

OpenStudy (zale101):

Step 1: Begin with the given grams of sulfuric acid, H2SO4 (which is 250 g ) Step 2: Multiply it by the molar mass of sulfuric acid. Why? Because we want to convert it to moles of sulfuric acid. That mean the conversion factor we are going to use for the molar mass is \(\frac{mols}{grams}\) so that the grams of the molar mass will cancel with 250 grams and leaves us with moles. In other words, you are dividing 250 by the molar mass of sulfuric acid.

OpenStudy (zale101):

However, we are not done, but before i continue, can you do these two steps?

OpenStudy (zale101):

If you have already done this question, please show your work so i can correct you if you have made any mistakes or if you have no absolute idea of where to begin then i suggest to start doing these two steps i provided above.

OpenStudy (18jonea):

98.079 x 250

OpenStudy (18jonea):

are those the right numbers to multiply

OpenStudy (zale101):

Please show your units, what are you canceling here?

OpenStudy (zale101):

@18jonea read what i wrote carefully for step two. I didn't mention multiplying.

OpenStudy (18jonea):

98.079g/mol x 250 g

OpenStudy (18jonea):

you said multiply it by the molecular masss

OpenStudy (18jonea):

oops sorry

OpenStudy (zale101):

@18jonea your goal is to cancel the grams.

OpenStudy (18jonea):

250 g / 98.079mol

OpenStudy (zale101):

write the appropriate conversion factor for the molar mass

OpenStudy (zale101):

Yes you divide

OpenStudy (18jonea):

2.549

OpenStudy (zale101):

But you have to make your units correct. \(250~grams \times \LARGE \frac{moles}{98.08~grams}\) so that the grams cancel out \(250~\cancel{grams }\times \LARGE \frac{moles}{98.08~\cancel{grams}}\)

OpenStudy (zale101):

what are the units for 2.549?

OpenStudy (18jonea):

mol?

OpenStudy (zale101):

Yes :)

OpenStudy (zale101):

Moles of sulfuric acid or H2SO4

OpenStudy (zale101):

So 2.549 mols of H2SO4

OpenStudy (zale101):

Now, your question is asking for the grams of sulfur in the sulfuric acid

OpenStudy (zale101):

In order to do that, you must know how to find the moles of sulfur in 2.549 sulfuric acid. .

OpenStudy (18jonea):

so how would i do that

OpenStudy (sweetburger):

*Edited post from above there was a typo You first need the molar mass of Sulfuric acid you then must divide the mass of sulfur by the total mass of Sulfuric acid this will give you the percent mass of Sulfur in Sulfuric acid. Then you take the 250 grams and multiply by the percent of sulfur found in sulfuric acid. This will give you the mass of sulfur in 250 grams of sulfuric acid.

OpenStudy (18jonea):

im so confused

OpenStudy (sweetburger):

mass of sulfur/ mass of sulfuric acid = ratio of sulfur in sulfuric acid multiply this value times the number of moles of Sulfuric acid that you found (2.549moles) this will give you the moles of Sulfur then you must multiply by the mass of Sulfur. It is honestly easier in this question to never convert to moles as it causes extra steps.

OpenStudy (sweetburger):

If your still confused let us know.

OpenStudy (18jonea):

ok can you help caue i honestly have no idea

OpenStudy (zale101):

how many atoms of sulfur is found in H2SO4? @18jonea

OpenStudy (18jonea):

1

OpenStudy (zale101):

Good, so what do you put for the mass of sulfur when calculating \(\Large \frac{mass~of sulfur}{mass~of~sulfuric~acid}\) @sweetburger suggested you to do.

OpenStudy (zale101):

Answer. Just the molar mass of Sulfur

OpenStudy (zale101):

1 mol of sulfur is what grams?

OpenStudy (zale101):

@18jonea

OpenStudy (18jonea):

34.0146

OpenStudy (zale101):

only 32.06 g/mol

OpenStudy (18jonea):

o ok

OpenStudy (zale101):

Google molar mass of sulfur and 32.06 g/mol will come up.

OpenStudy (sweetburger):

Are we allowed to post links?

OpenStudy (18jonea):

i just did that

OpenStudy (zale101):

Yes. Links for educational purposes

OpenStudy (sweetburger):

http://www.ptable.com/ pretty cool interactive periodic table that you can leave on your favorites bar

OpenStudy (zale101):

@sweetburger that's awesome!

OpenStudy (zale101):

@18jonea if you have a periodic table, you can always look at the molar mass of elements. They are all there.

OpenStudy (zale101):

@18jonea please show your ratio calculations when doing \(\Large\frac{ mass~sulfur}{mass~sulfuric~acid}\)

OpenStudy (zale101):

Can you do this now?

OpenStudy (18jonea):

32.06/2.549

OpenStudy (zale101):

The denominator is incorrect, it's asking for the molar mass of sulfuric acid. Actually, let me change the wording. \(\Large\frac{ Molar~~mass~of~sulfur}{Molar~mass~of~sulfuric~acid}\)

OpenStudy (18jonea):

98.079

OpenStudy (18jonea):

is that right?

OpenStudy (18jonea):

.33 g

OpenStudy (zale101):

Yes. That is the molar mass of H2SO4

OpenStudy (zale101):

Correct, 0.33 is the ratio

OpenStudy (zale101):

Now read the instructions what @sweetburger said after solving for the ratio.

OpenStudy (zale101):

What does it say?

OpenStudy (18jonea):

ratio of sulfur in sulfuric acid multiply this value times the number of moles of Sulfuric acid that you found (2.549moles) this will give you the moles of Sulfur then you must multiply by the mass of Sulfur.

OpenStudy (zale101):

Do as he says.

OpenStudy (18jonea):

.33 x 2.549= .84117

OpenStudy (zale101):

Correct.

OpenStudy (zale101):

What are the units of 0.841?

OpenStudy (18jonea):

g?

OpenStudy (zale101):

No. You multiplied the ratio with 2.549 moles of h2so4. Correct?

OpenStudy (zale101):

How did the answer turn into grams..?

OpenStudy (18jonea):

so moles

OpenStudy (zale101):

That's right, but moles of what?

OpenStudy (18jonea):

i dont know

OpenStudy (18jonea):

?

OpenStudy (zale101):

Moles of Sulfur.

OpenStudy (18jonea):

ok so now what do i do

OpenStudy (zale101):

The last step is to use the moles of sulfur and multiply by the molar mass of sulfur (g/mol) so that the moles cancel out and you are left with just grams of sulfur.

OpenStudy (18jonea):

so what would that be?

OpenStudy (18jonea):

?

OpenStudy (zale101):

I want you to do.

OpenStudy (18jonea):

what are the numbers

OpenStudy (18jonea):

would it be .841x 32.06= 26.96

OpenStudy (18jonea):

@Zale101

OpenStudy (18jonea):

@sweetburger

OpenStudy (18jonea):

@Zale101

OpenStudy (18jonea):

but those arent any of the answers

OpenStudy (zale101):

what is the answer or choices given?

OpenStudy (18jonea):

16 g 0.33g 160 g 82 g

OpenStudy (zale101):

Oh I see. I think you need to multiply the ratio or percent of sulfur found in H2SO4 by the given mass of H2SO4 (250 grams). You'll get something around 82 grams That's all is to it...

OpenStudy (18jonea):

so the answer is 82 g

OpenStudy (18jonea):

?

OpenStudy (zale101):

Another way to get the answer is to use the moles of H2SO4 we solved for (2.549) and use mole to mole ratio. \(2.549~moles~h2So4 \times \Large \frac{1~mole~S~in h2so4}{1~moles~h2so4}=2.549~moles ~of~Sulfur \) Then you multiply the moles of sulfur (2.549 mol) by the molar mass of sulfur (32 grams per mole and you will get 82 grams of sulfur.